# Improving Cramér–Rao Bound And Its Variants: An Extrinsic Geometry Perspective

### v1

@@ 5 Analytical examples
It is easy to see that $\mathcal{B}^{(2)}_{\mathrm{jet}}=\mathcal{C}^{(1)}$ always; that is, this is not example dependent. Therefore, also in our example,
$\mathcal{B}^{(2)}_{\mathrm{jet}}\;=\;\|\operatorname{Proj}_{\mathcal{T}_{2}}(H_{3}s_{0})\|^{2}\;=\;\frac{432}{\,2c^{2}+4c+173\,}.$
In particular, $\mathcal{B}^{(2)}_{\mathrm{jet}}>0$ for every real $c$. Since $\mathcal{B}^{(2)}_{\text{score}}=0$ and $\mathcal{C}^{(2)}\geq\mathcal{B}^{(2)}_{\mathrm{jet}}>0$ for all $c\in\mathbb{R}$ (of course, we are concerned with $c\neq 0$ in this example), we have demonstrated strict improvement also at $m=2$ with Theorem 5.
Explicitly, we have
$\eta_{3}/s_{0}=(H_{1}+H_{2})^{3}+3c\,(H_{1}H_{2}+H_{2}^{2})=H_{6}+3H_{5}+(15+3c)H_{4}+(25+3c)H_{3}$
$(45+12c)H_{2}+(33+6c)H_{1}+(14+6c)H_{0}.$
The tangent space at order $m=2$ is $\mathcal{T}_{2}=\operatorname{span}\{\eta_{1},\eta_{2}\}$, and the second fundamental vector is
$\mathrm{II}_{2}\;=\;\eta_{3}-\operatorname{Proj}_{\mathcal{T}_{2}}\eta_{3}.$
The residual variance bound at order $m=2$ is then given by
$\mathcal{C}^{(2)}=\big\|\operatorname{Proj}_{\mathcal{T}_{2}}(H_{3}s_{0})\big\|^{2}+\big\|\operatorname{Proj}_{\operatorname{span}\{\mathrm{II}_{2}\}}(H_{3}s_{0})\big\|^{2}.$
Recall
$\mathcal{C}^{(1)}=\frac{432}{\,2c^{2}+4c+173\,}$,
and a direct symbolic computation yields
$\mathcal{C}^{(2)}-\mathcal{C}^{(1)}\;=\;\frac{18\bigl(6c^{3}+38c^{2}-209c+65\bigr)^{2}}{(2c^{2}+4c+173)\,\bigl(306c^{4}+2328c^{3}+9349c^{2}+16070c+216283\bigr)}.$
Substituting $c=1$, we obtain
$\mathcal{C}^{(2)}-\mathcal{C}^{(1)}\approx 0.0041.$
The improvement is minimal for this particular value of $c$. Can we better this significantly using the stronger Proposition 8?
@@
$\eta_{2}\;=\;\tfrac{1}{2}\widetilde{Y}_{2}+W_{2},\qquad W_{2}\;=\;\tfrac{1}{4}Y_{1}^{2}s_{0}=s_{0}\big(2+6H_{1}+4H_{2}+3H_{3}+H_{4}\big).$
@@
$\|W_{2}\|^{2}$ $=2^{2}\cdot 0!+6^{2}\cdot 1!+4^{2}\cdot 2!+3^{2}\cdot 3!+1^{2}\cdot 4!$
$=4+36+32+54+24=150,$
$\|\eta_{1}\|^{2}$ $=1^{2}\cdot 1!+1^{2}\cdot 2!=1+2=3,$
$\|\eta_{2}\|^{2}$ $=3^{2}\cdot 0!+4^{2}\cdot 1!+(5+c)^{2}\cdot 2!+2^{2}\cdot 3!+1^{2}\cdot 4!$
$=9+16+2(5+c)^{2}+24+24=2c^{2}+20c+123,$
@@
$\langle W_{2},\eta_{1}\rangle$ $=6\cdot 1!+4\cdot 2!=6+8=14,$
$\langle W_{2},\eta_{2}\rangle$ $=2\cdot 3\cdot 0!+6\cdot 4\cdot 1!+4(5+c)\cdot 2!+3\cdot 2\cdot 3!+1\cdot 1\cdot 4!$
$=6+24+8(5+c)+36+24=130+8c.$
@@
i.e.
$\begin{pmatrix}3&14+2c\\[4.0pt] 14+2c&2c^{2}+20c+123\end{pmatrix}\begin{pmatrix}a\\ b\end{pmatrix}=\begin{pmatrix}14\\ 130+8c\end{pmatrix}$.
One obtains, after simplification, the closed form
$W_{2}^{\perp}\;=\;s_{0}\Bigg[\frac{4\big(c^{2}+5c-59\big)}{\,2c^{2}+4c+173\,}\,H_{0}\;+\;\frac{12\big(11c+30\big)}{\,2c^{2}+4c+173\,}\,H_{1}\;-\;\frac{6\big(11c+30\big)}{\,2c^{2}+4c+173\,}\,H_{2}\;$
$\frac{6c^{2}+20c+131}{\,2c^{2}+4c+173\,}\,H_{3}\;+\;\frac{2c^{2}+8c-21}{\,2c^{2}+4c+173\,}\,H_{4}\Bigg],$
and
$\|W_{2}^{\perp}\|^{2}\;=\;\frac{2\big(82c^{2}+428c+1051\big)}{\,2c^{2}+4c+173\,}>0.$
@@
We note that the coefficient of $H_{3}$ in $W_{2}^{\perp}$, which is $\frac{6c^{2}+20c+131}{\,2c^{2}+4c+173\,}$, is not zero for any $c\in\mathbb{R}$. With $\mathcal{N}_{2}^{\mathrm{aug}}=\mathrm{span}\{\Pi_{2},W_{2}^{\perp}\}$ and $\widetilde{\mathcal{T}}_{2}=\mathrm{span}\{H_{1}s_{0},H_{2}s_{0}\}$, this implies that the sufficient conditions $\operatorname{Proj}_{\mathcal{N}_{2}^{\mathrm{aug}}}H_{3}s_{0}\neq 0$ and $\mathcal{N}_{2}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{2}$ are satisfied, and Proposition 8 guarantees that
@@
Using explicit computations, we have
$\mathcal{C}^{(2)}_{\mathrm{aug}}\;\geq\;\big\|\operatorname{Proj}_{\mathcal{T}_{2}}(H_{3}s_{0})\big\|^{2}\;+\;\frac{\langle H_{3}s_{0},W_{2}^{\perp}\rangle^{2}}{\|W_{2}^{\perp}\|^{2}}\;=\;\frac{18\,(18c^{2}+84c+245)}{\,82c^{2}+428c+1051\,}$,
which approximately evaluates to $4.001$ at $c=1$, significantly improving over $\mathcal{C}^{(1)}$ and $\mathcal{C}^{(2)}$ (both of which were approximately $2.41$); in fact, one can establish that $\mathcal{C}^{(2)}_{\mathrm{aug}}>\mathcal{C}^{(1)}$ for all $c$. Note that the lower bound in the last display is due to the fact that we have not included $\Pi_{2}$ in computing the second term and, in fact, $\mathcal{N}_{2}^{\mathrm{aug}}=\mathrm{span}\{\Pi_{2},W_{2}^{\perp}\}$.
@@ 6 Conclusion
In this work we developed refinements of the classical Cramér-Rao lower bound, and some of its own refinements such as the Bhattacharyya-type bounds that use higher-order score functions, in a manner consistent with the local differential geometry of the statistical model. The CRB case incorporated curvature-aware corrections derived from the second fundamental form on tangent vectors of the statistical manifold embedded into a fixed Hilbert space using the square root embedding, while the higher-order version also considered the decomposition of the corresponding jets in terms of the true scores using a special case of the Faà di Bruno formula. We detailed explicit examples where the geometric correction terms result in definite improvements over the CRB and its variants. This confirms that curvature-aware corrections, when properly projected and interpreted within a Riemannian framework, yield provable and practically relevant enhancements. Further research may examine the generality and limitations of these geometric corrections in broader model classes, such as those with higher-order nonlinearity in the statistical embedding or with weaker assumptions on regularity than what we allowed, for instance. The connection between extrinsic curvature and information-theoretic optimality, particularly beyond the second-order regime, remains an area with substantial theoretical and methodological potential.
@@ Acknowledgements.
We are thankful to Prof. Pillai for sharing his paper [11], which formed the motivation for this work.

### v2

@@ 5 Analytical examples
It is immediate that
$\mathcal{B}^{(2)}_{\mathrm{jet}}\;=\;\|\operatorname{Proj}_{\mathcal{T}_{2}}(H_{3}s_{0})\|^{2}\;=\;\mathcal{C}^{(1)}.$
In particular, $\mathcal{B}^{(2)}_{\mathrm{jet}}>0$ for every real $c$. Since $\mathcal{B}^{(2)}_{\text{score}}=0$ and $\mathcal{C}^{(2)}\geq\mathcal{B}^{(2)}_{\mathrm{jet}}>0$ for all $c\in\mathbb{R}$, we have demonstrated strict improvement also at $m=2$ with Theorem 5.
@@
$\eta_{2}\;=\;\tfrac{1}{2}\widetilde{Y}_{2}+W_{2},\qquad W_{2}\;=\;\tfrac{1}{4}Y_{1}^{2}s_{0}=s_{0}\big(H_{4}+2H_{3}+5H_{2}+4H_{1}+3H_{0}\big).$
@@
$\|W_{2}\|^{2}$ $=123$
$\|\eta_{1}\|^{2}$ $=3,$
$\|\eta_{2}\|^{2}$ $=2c^{2}+20c+123,$
@@
$\langle W_{2},\eta_{1}\rangle$ $=14,$
$\langle W_{2},\eta_{2}\rangle$ $=105+10c.$
@@
Solving, we obtain the explicit coefficients
$a=\frac{8c^{2}-70c+252}{2c^{2}+4c+173},\qquad b=\frac{2c+119}{2c^{2}+4c+173}$
The residual (normal) component is
$W_{2}^{\perp}=W_{2}-a\,\eta_{1}-b\,\eta_{2}.$
Using
$\|W_{2}^{\perp}\|^{2}=\|W_{2}\|^{2}-\big(a\langle W_{2},\eta_{1}\rangle+b\langle W_{2},\eta_{2}\rangle\big),$
together with $\|W_{2}\|^{2}=123$, we find
$\|W_{2}^{\perp}\|^{2}=\frac{114c^{2}+72c+5256}{2c^{2}+4c+173}.$
Since the $H_{3}$ coefficient of $W_{2}^{\perp}$ is
22b=2(1b),2-2b=2(1-b),
we obtain
$\langle H_{3}s_{0},W_{2}^{\perp}\rangle=12(1-b),\qquad\langle H_{3}s_{0},W_{2}^{\perp}\rangle^{2}=144(1-b)^{2}.$
Figure 1: The ratio $\frac{\langle H_{3}s_{0},W_{2}^{\perp}\rangle^{2}}{\|W_{2}^{\perp}\|^{2}}$ as a function of the parameter $c$.
Therefore,
$\frac{\langle H_{3}s_{0},W_{2}^{\perp}\rangle^{2}}{\|W_{2}^{\perp}\|^{2}}=\frac{96(c^{2}+c+27)^{2}}{(2c^{2}+4c+173)(19c^{2}+12c+876)}.$
@@
Note that the residual $W_{2}^{\perp}$ has a non-vanishing component in the $H_{3}$ direction for every $c\in\mathbb{R}$. With $\mathcal{N}_{2}^{\mathrm{aug}}=\mathrm{span}\{W_{2}^{\perp}\}$ and $\widetilde{\mathcal{T}}_{2}=\operatorname{span}\{\widetilde{Y}_{1}(0),\widetilde{Y}_{2}(0)\}=\operatorname{span}\{2(H_{1}+H_{2})s_{0},\;(2c\,H_{2}-12H_{0})s_{0}\}$, this implies that the sufficient conditions $\operatorname{Proj}_{\mathcal{N}_{2}^{\mathrm{aug}}}(H_{3}s_{0})\neq 0$ and $\mathcal{N}_{2}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{2}$ are satisfied for all real $c$, and Proposition 8 guarantees that $\mathcal{C}_{\mathrm{aug}}^{(2)}\;>\;\mathcal{B}_{\mathrm{score}}^{(2)},$ with the correction along the new normal direction in the augmented space as computed explicitly above. For reader convenience, we provide a visualization of the correction we get by means of a graph that depicts how $\frac{\langle H_{3}s_{0},W_{2}^{\perp}\rangle^{2}}{\|W_{2}^{\perp}\|^{2}}$ varies with $c$ in Figure 1.
Finally, we consider a variation of Example 1 to validate our claims in Proposition 8 regarding possible improvements over $\mathcal{B}_{\mathrm{score}}^{(m)}$ by explicitly using the extrinsic directions $W_{2}^{\perp},\dots,W_{m}^{\perp}$ in addition to $\mathcal{B}^{(2)}_{\text{jet}}$, serving to compare lower bounds obtained using only the $\{\eta_{k}\}_{k=1}^{m}$ and those that use $\{\widetilde{Y}_{k}\}_{k=1}^{m}$.
Example 4 (Strict improvement of $\mathcal{C}_{\mathrm{aug}}^{(2)}$ over $\mathcal{B}_{\mathrm{score}}^{(2)}$).
Consider the one-dimensional location model
$f(x;\theta)=\frac{1}{Z}\exp\!\left(-(x-\theta)^{4}-\alpha(x-\theta)^{3}\right),\qquad\alpha\neq 0,$
where
$Z=\int_{\mathbb{R}}\exp(-u^{4}-\alpha u^{3})\,du<\infty,\quad u:=x-\theta.$
This model is smooth, regular, non-symmetric, and lies outside the exponential family.
Unbiased estimator.
Since this is a location family, define
$\mu:=\mathbb{E}_{0}[u]=\frac{1}{Z}\int_{\mathbb{R}}u\,e^{-u^{4}-\alpha u^{3}}\,du,$
and consider the estimator
$T(X):=X-\mu.$
Then
$\mathbb{E}_{\theta}[T(X)]=\theta+\mathbb{E}_{0}[u]-\mu=\theta,$
so $T$ is unbiased for all $\theta$. The centered error is
$Z_{0}:=T(X)-\theta=u-\mu.$
Scores and score images.
The log-density is
$\log f(x;\theta)=-u^{4}-\alpha u^{3}-\log Z,$
hence the first two score functions are
$Y_{1}(u)=4u^{3}+3\alpha u^{2},\qquad Y_{2}(u)=-12u^{2}-6\alpha u.$
Let $s_{\theta}=\sqrt{f(\cdot;\theta)}$ and $\widetilde{Y}_{j}=Y_{j}s_{\theta}$.
Jets via Bell polynomials.
Using the Bell polynomial identity
$\eta_{k}=\partial_{\theta}^{k}s_{\theta}=s_{\theta}\,B_{k}\!\Big(\tfrac{1}{2}Y_{1},\dots,\tfrac{1}{2}Y_{k}\Big),$
the first two jets are
$\eta_{1}$ $=\tfrac{1}{2}Y_{1}s_{\theta}=s_{\theta}\Big(2u^{3}+\tfrac{3}{2}\alpha u^{2}\Big),$
$\eta_{2}$ $=s_{\theta}\Big(\tfrac{1}{4}Y_{1}^{2}+\tfrac{1}{2}Y_{2}\Big).$
Since
$Y_{1}^{2}=16u^{6}+24\alpha u^{5}+9\alpha^{2}u^{4},$
we obtain
$\eta_{2}=s_{\theta}\Big(4u^{6}+6\alpha u^{5}+\tfrac{9}{4}\alpha^{2}u^{4}-6u^{2}-3\alpha u\Big).$
Bell remainder and augmented normal direction.
By definition,
$\eta_{2}=\tfrac{1}{2}Y_{2}s_{\theta}+W_{2},$
where the Bell remainder is
$W_{2}=s_{\theta}\Big(4u^{6}+6\alpha u^{5}+\tfrac{9}{4}\alpha^{2}u^{4}\Big).$
Let
$\mathcal{T}_{2}=\operatorname{span}\{\eta_{1},\eta_{2}\},\qquad\mathcal{N}_{2}^{\mathrm{aug}}=\operatorname{span}\{W_{2}^{\perp}\}\subset\mathcal{T}_{2}^{\perp}.$
Strict improvement at order $m=2$.
The estimator residual in $L^{2}(\mu)$ is
$\widetilde{Z}_{0}=(u-\mu)s_{\theta},$
which has a nonzero odd component. Notice that $W_{2}$ contains the odd term $6\alpha u^{5}s_{\theta}$ and $\alpha\neq 0$.
The second-order score-based Bhattacharyya bound is
$\mathcal{B}_{\mathrm{score}}^{(2)}=\big\|\operatorname{Proj}_{\widetilde{\mathcal{T}}_{2}}\widetilde{Z}_{0}\big\|^{2},\qquad\widetilde{\mathcal{T}}_{2}=\operatorname{span}\{\widetilde{Y}_{1},\widetilde{Y}_{2}\}.$
The augmented bound is
$\mathcal{C}_{\mathrm{aug}}^{(2)}=\big\|\operatorname{Proj}_{\mathcal{T}_{2}}\widetilde{Z}_{0}\big\|^{2}+\big\|\operatorname{Proj}_{\mathcal{N}_{2}^{\mathrm{aug}}}\widetilde{Z}_{0}\big\|^{2}.$
For generic asymmetric quartic models, the conditions $\operatorname{Proj}_{\mathcal{N}_{2}^{\mathrm{aug}}}\widetilde{Z}_{0}\neq 0$ and $\mathcal{N}_{2}^{\mathrm{aug}}\not\subseteq\widetilde{\mathcal{T}}_{2}$ are satisfied, and Proposition 8 implies
@@
The magnitude of the improvement obviously depends on the asymmetry parameter $\alpha$. For instance, with $\alpha=1.95$, we get the numerically verifiable values $\mathcal{B}_{\mathrm{score}}^{(2)}\approx 0.16,\quad C_{\mathrm{aug}}^{(2)}\approx 0.19$. This strict improvement occurs at fixed order $m=2$ and is due, in this particular example, mainly to the $\mathcal{B}_{\mathrm{jet}}^{(2)}$ component in $\mathcal{C}_{\mathrm{aug}}^{(2)}$.
Remark 5.
Our primary goal in these examples is conceptual: to demonstrate that statistical inefficiency can be interpreted geometrically as curvature of the embedded statistical manifold in $L^{2}(\mu)$. The Hermite–Gaussian examples are deliberately chosen for analytical transparency, allowing all geometric quantities (jets, second fundamental form, normal components) to be computed explicitly and rigorously. While the resulting correction depends on the estimator and may not always be computationally convenient, this dependence is intrinsic: inefficiency itself is estimator-specific. Importantly, the framework applies beyond toy examples. For instance, constrained or misspecified estimators and semiparametric models with nuisance parameters could give rise to nonzero curvature terms and admit strict refinements over the CRB. Furthermore, we also demonstrated a possible numerical computation of our bounds using Example 4 where, in fact, our primary interest was to illustrate Proposition 8.
@@ 6 Conclusion
In this work we developed (non-asymptotic) refinements of the classical Cramér-Rao lower bound, and some of its own refinements such as the Bhattacharyya-type bounds that use higher-order score functions, in a manner consistent with the local differential geometry of the statistical model. The CRB case incorporated curvature-aware corrections derived from the second fundamental form on tangent vectors of the statistical manifold embedded into a fixed Hilbert space using the square root map, while the higher-order version also relied on a natural extension. Furthermore, we considered the decomposition of the corresponding jets in terms of the true scores using a special case of the Faà di Bruno formula to demonstrate improvements over Bhattacharyya-type bounds that use log-likelihood derivatives. We detailed explicit examples where the geometric correction terms result in definite improvements over the CRB and its variants. This confirms that curvature-aware corrections, when properly projected and interpreted within a Riemannian framework, yield provable and relevant inefficiency measures. Further research may examine the generality and limitations of these geometric corrections in broader model classes, such as those with higher-order nonlinearity in the statistical embedding or with weaker assumptions on regularity than what we allowed, for instance. The connection between extrinsic curvature and information-theoretic optimality, particularly beyond the second-order regime, remains an area with substantial theoretical and methodological potential.
@@ Acknowledgements.
We are thankful to Prof. Pillai for sharing his article [15], which formed the initial motivation for this work.